OpenStudy (anonymous):

Find f. f ''(t) = 7et + 9 sin t, f(0) = 0, f(π) = 0

OpenStudy (anonymous):

7e^t

OpenStudy (sidsiddhartha):

ok then start with integrating

OpenStudy (anonymous):

yeah im having trouble with the last step

OpenStudy (anonymous):

plugging in pi

OpenStudy (sidsiddhartha):

ok i,m doing it verify it ok?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

f(t) = 7e^t-9sint+Ct+D

OpenStudy (sidsiddhartha):

\[f''(t)=7e^t+9sint\\f'(t)=7e^t-9cost+c\] now find f'(0)

OpenStudy (anonymous):

its f(0) so we dont plug it in yet until we integrate it one more time right?

OpenStudy (paxpolaris):

\[f(t) = 7e^t-9\sin t+C t +D\] so you plug in f(0)=0 ... what do you get for D

OpenStudy (anonymous):

-7

OpenStudy (sidsiddhartha):

now put D=-7 and f(pi)=0

OpenStudy (paxpolaris):

\[f(t) = 7e^t-9\sin t+C t -7\] and plugging \(f(\pi)=0\) ...

OpenStudy (anonymous):

-((7e^pi)/pi) +7

OpenStudy (paxpolaris):

\[0 =7e^\pi+\pi C -7\] \[\large \implies C= {7-7e^\pi \over \pi}\]

OpenStudy (anonymous):

thank you!